OpenStudy (anonymous):

f(x)=-2x^3-x^2+4x+4 Find the average slope of this function on the interval(1,2)= By the mean value theorem, we know there exists a c in the open interval(1,2) such that f'(c) is equal to this mean slope. Find the value of c in the interval which works..

OpenStudy (anonymous):

you need two numbers first \[f(2)\] and \[f(1)\] did you find them?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what are they?

OpenStudy (anonymous):

-8 and 5

OpenStudy (anonymous):

ok then \[\frac{f(2)-f(1)}{2-1}=\frac{-8-5}{1}=-13\]

OpenStudy (anonymous):

next you need the derivative what is that?

OpenStudy (anonymous):

-6x^2-2x+4

OpenStudy (anonymous):

ok final job is to solve \[-6x^2-2x+4=-8\] and you are done

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wyhooo mistakE!!

OpenStudy (anonymous):

\[-6x^2-2x+4=-13\] is more like it

OpenStudy (jhannybean):

About to say xD

OpenStudy (anonymous):

hmm ok

OpenStudy (anonymous):

yeah must be my bedtime